OpenStudy (anonymous):

I am so very lost. Can Someone help me solve this? A test charge of -5.0 x 10-7 C is placed between other charges so that it is 50 mm from a charge of -3.0 x 10-7 C and 10 cm from a charge of -6.0 x 10-7 C. The three charges lie along a straight line. The answer is 0.27N to the right. Idk how to get that?

OpenStudy (aaronq):

So you know coulomb's law, \(F=k\dfrac{q_1q_2}{r^2}\)?

OpenStudy (aaronq):



OpenStudy (anonymous):

it's 5mm though.

OpenStudy (aaronq):

it says 50 mm which is 5 cm

OpenStudy (anonymous):

oh. Okay. I tried using that formula, I wrote Fnet = f1 + f2 and then tried solving for both, but I didn't get the right answer

OpenStudy (aaronq):

yeah, that should be right. Did you use SI units, i.e. convert distance to meters?

OpenStudy (anonymous):

Yes.

OpenStudy (aaronq):

hm gimme me a second, i'm going to try it

OpenStudy (aaronq):

i got 0.2696265 N which is 0.27 N I think the problem you're having the signs it should look like this (notice one is positive and one negative) \(\large F=kq_1[\dfrac{q_2}{r^2}+-\dfrac{q_3}{r^2}]\) because both charges are putting a repulsive force on the test charge (all are negative) 

OpenStudy (anonymous):

How do you know that one is negative and not both?

OpenStudy (aaronq):

because of the signs of the charges and their interaction - all are negative, so they are repelling on another Along the axis, each one is exerting a force, this is a vector and thus according to the axis is set it up they are opposing forces

OpenStudy (anonymous):

Okay, I understand how they interact. I'm just uncertain how you would determine what the charge is. Sorry

OpenStudy (aaronq):

it's given in the question -5.0 x 10-7 C notice the negative sign, that is the sign of the chargw

OpenStudy (anonymous):

Oh, my physics teacher told me that it wasn't

OpenStudy (anonymous):

Thank you so much for the help. Things are so much clearer now.

OpenStudy (aaronq):

no problem! glad i could be of assistance